Solve with complete steps to get 100% feedback!!
15 4. Find the solution of the initial value problem y + y - 2y = 2t, y(0) = 0, y(0) = 1
Answer #1

CO) = 0, we know, y+y-2y = 22 - y y Con=1 Taking Laplace Transform, dry s F(s)- s fco) -fco) – at dy sfcs) - fco) - dt Sub2ts² + B A ++D S²(S-1)(3+2) (S-1) (5+2) S Alin 2+5² = A(G²) (5+2) + B (5²) (S-1)+(2+D put s=1 3= 3A + CAD put S = -2 3= 3A+C+

For any specific explanation please drop a comment.